# A Superconducting Single-Atom Phonon Laser

### v2

@@ III.1 Considerations of Qubit Anharmonicity
We modeled the single-atom laser using a two-level approximation; however, under strong drives, it is important to consider the higher levels of the transmon qubit. We can instead model the qubit as a Kerr oscillator with a large anharmonicity $\alpha$. Therefore, we can re-write the Hamiltonian in the form
@@
Where we have replaced the two-level raising and lowering operators with bosonic raising and lowering operators $\hat{c}^{(\dagger)}$. If we calculate the number of steady-state phonons comparing the Kerr oscillator model to the two-level system model, we find at powers below and near the upper lasing threshold, the population is slightly larger in the Kerr oscillator model; see Fig. S8. There exists a discrepancy at high drive powers; however, this was not observed within the experiment since we likely could not apply a large enough drive without heating the sample. Thus, we are confident that modeling our experiment as a two-level system is accurate, especially for calculating the upper lasing threshold. Moreover, this was required for numerical efficiency. The two-level system simulation in Fig. S8 ran for approximately 20 minutes, whereas considering only three levels of the Kerr oscillator increased the simulation time to approximately 10 hours.
Figure S8: Two-Level verses Anharmonic Oscillator Simulations. (A) The simulated phonon population when considering the transmon qubit a two-level system versus considering three levels of the Kerr oscillator.
For numerical efficiency, we have constructed a simplified model where we have removed the coupling to the readout resonator. Studying Eq. S3, the readout resonator broadens the qubit and causes a constant frequency offset due to steady-state photons in the cavity. In this section, we aim to explore the properties of the single-atom laser. The Hamiltonian of our simplified model is given by
@@
and we choose to have the qubit and phonon resonant for the simulation. Moreover, we have increased intrinsic qubit linewidth $\Gamma_{\rm q}=1.5$ MHz to account for the lack of readout resonator broadening. We also increased the phonon linewidth $\gamma_{\rm b}=25$ kHz to reduce the size of the Hilbert space for numerical efficiency.
@@ III.2 Phonon Phase Coherence
First, we aim to investigate if the drive tone seeds the phase of the phonon mode. In the model described in Ref. Ashhab et al. (2009) the phase of the lasing mode is random and results in a distribution in the IQ plane. However, if we examine the phase space distribution of the phonon mode in our model, we find that the phase of the drive tone instead sets the phase; see Fig. S9. This results from the weak hybridization between the qubit and phonon mode. When the drive is switched on, the weak hybridization enables the seeding of the phase of the phonon mode prior to the onset of lasing. This is followed by the stimulated emission of phonons, which produces a large coherent state with a set phase. The ability to set the phase of the phonon mode is important for applications such as the generation of Schrödinger cat states.
Figure S9: Phonon Lasing Phase Seeding. (A) Phonon population as a function of qubit drive power. Colored points correspond to the Wigner plots (B-E). (B-E) Wigner distribution plots corresponding to the points in (A). The phase of the lasing state is set by the drive phase; here, the drive phase was $\pi/4$. The plots are shown starting from the lowest power (B) to the largest power (E).
@@ III.3 Phonon Mode Anharmonicity
Finally, the coupling between the qubit and phonon mode raises the question of how much anharmonicity is inherited by the phonon mode. In the strong coupling limit, the phonon mode will inherit half of the qubit anharmonicity and, thus, would be unable to lase since its Hilbert space is truncated to a single excitation. From the transparency window measurement, we observe little frequency shift of the phonon mode, suggesting the mode is well approximated as a harmonic oscillator and not a Duffing oscillator. However, again, we can utilize our simplified model to explore the inherited anharmonicity by directly driving the phonon mode in the simulation, bypassing all lasing dynamics. We find that at a drive strength that generates, on average, fifty phonons, similar to what was observed in the experiment, the phonon mode has a negligible frequency shift. Suggesting that in the weak hybridization limit, the phonon mode can be treated to a high degree as a linear harmonic oscillator.

### v3

@@ III.1 Considerations of Qubit Anharmonicity
We modeled our system using a two-level approximation; however, under strong drives, it is important to consider the higher levels of the transmon qubit. We can instead model the qubit as a Kerr oscillator with a large anharmonicity $\alpha$. Therefore, we can re-write the Hamiltonian in the form
@@
Where we have replaced the two-level raising and lowering operators with bosonic raising and lowering operators $\hat{c}^{(\dagger)}$. If we calculate the number of steady-state phonons comparing the Kerr oscillator model to the two-level system model, we find that at powers below and near the upper threshold, the population is slightly larger in the Kerr oscillator model; see Supplementary Fig. 9. There exists a discrepancy at high drive powers; however, this was not observed within the experiment since we likely could not apply a large enough drive without heating the sample. Thus, we are confident that modeling our experiment as a two-level system is accurate, especially for calculating the upper threshold. Moreover, this was required for numerical efficiency. The two-level system simulation in Supplementary Fig. 9 ran for approximately 20 minutes, whereas considering only three levels of the Kerr oscillator increased the simulation time to approximately 10 hours.
SUPPLEMENTARY FIGURE 9: Two-Level versus Anharmonic Oscillator Simulations. The simulated phonon population when considering the transmon qubit as a two-level system versus considering three levels of the Kerr oscillator.
For numerical efficiency, we have constructed a simplified model where we have removed the coupling to the readout resonator. Studying Eq. S3, the readout resonator broadens the qubit and causes a constant frequency offset due to steady-state photons in the cavity. The Hamiltonian of our simplified model is given by
@@
and we choose to have the qubit and phonon resonant for the simulation. Moreover, we have increased the intrinsic qubit linewidth $\Gamma_{\rm q}/2\pi=1.5$ MHz to account for the lack of readout resonator broadening. We also increased the phonon linewidth $\gamma_{\rm b}/2\pi=25$ kHz to reduce the size of the Hilbert space for numerical efficiency.
@@ III.2 Phonon Phase Coherence
First, we aim to investigate if the drive tone seeds the phase of the phonon mode. In the model described in Ref. Ashhab et al. (2009) the phase of the phonon mode is random and results in a distribution in the IQ plane. However, if we examine the phase space distribution of the phonon mode in our model, we find that the phase of the drive tone instead sets the phase; see Supplementary Fig. 10. This results from the weak hybridization between the qubit and phonon mode. When the drive is switched on, the weak hybridization enables the seeding of the phase of the phonon mode. This is followed by the stimulated emission of phonons, which produces a large coherent state with a set phase. The ability to set the phase of the phonon mode is important for applications such as the generation of Schrödinger cat states.
SUPPLEMENTARY FIGURE 10: Phonon Phase Seeding. (A) Phonon population as a function of qubit drive power. Colored points correspond to the Wigner plots (B-E). (B-E) Wigner distribution plots corresponding to the points in (A). The phase of the coherent state is set by the drive phase; here, the drive phase was $\pi/4$. The plots are shown starting from the lowest power (B) to the largest power (E).
@@ III.3 Phonon Mode Anharmonicity
Finally, the coupling between the qubit and phonon mode raises the question of how much anharmonicity is inherited by the phonon mode. In the strong coupling limit, the phonon mode will inherit half of the qubit anharmonicity and, thus, would be unable to populate the phonon mode since its Hilbert space is truncated to a single excitation. From the transparency window measurement, we observe little frequency shift of the phonon mode, suggesting the mode is well approximated as a harmonic oscillator and not a Duffing oscillator. However, again, we can utilize our simplified model to explore the inherited anharmonicity by directly driving the phonon mode in the simulation. We find that at a drive strength that generates, on average, fifty phonons, similar to what was observed in the experiment, the phonon mode has a negligible frequency shift. Suggesting that in the weak hybridization limit, the phonon mode can be treated to a high degree as a linear harmonic oscillator.